# Quality-Diversity Generative Sampling for Learning with Synthetic Data

### 22 Dec 2023 v1

@@ Abstract
Generative models can serve as surrogates for some real data sources by creating synthetic training datasets, but in doing so they may transfer biases to downstream tasks. We focus on protecting quality and diversity when generating synthetic training datasets. We propose quality-diversity generative sampling (QDGS), a framework for sampling data uniformly across a user-defined measure space, despite the data coming from a biased generator. QDGS is a model-agnostic framework that uses prompt guidance to optimize a quality objective across measures of diversity for synthetically generated data, without fine-tuning the generative model. Using balanced synthetic datasets generated by QDGS, we first debias classifiers trained on color-biased shape datasets as a proof-of-concept. By applying QDGS to facial data synthesis, we prompt for desired semantic concepts, such as skin tone and age, to create an intersectional dataset with a combined blend of visual features. Leveraging this balanced data for training classifiers improves fairness while maintaining accuracy on facial recognition benchmarks. Code available at: github.com/Cylumn/qd-generative-sampling
Keywords: Balancing Datasets; Bias, Fairness & Privacy
@@ 1 Introduction
Balancing synthetic datasets is not a trivial task, since generative models harbor biases that relate to different semantic concepts. Deep generative models are often trained to replicate Internet data, where there are common issues of sampling selection bias. Internet data can also contain social biases that result from disparities in language use among social groups, eliciting stereotypes and prejudices (Otterbacher 2018). Biases may also be affected in training, through data augmentation and regularization methods that amplify class imbalance (Balestriero, Bottou, and LeCun 2022). While it is intractable to remove disparities across all semantic concepts, it is helpful to consider more than one at a time when creating synthetic datasets.
@@ 6 Results and Discussion
6.1 Spread of generative data distributions
@@
Table 3: Ablation results: QDGS is helpful for either loss.
@@
[…]
@@ B.2 Measure language prompts
  • $\bm{x}_{m1}=$“{red, blue}.” $\bm{x}_{m2}=$A {square, triangle}.”
  • $\bm{x}_{m1}=$A {red, blue} shape.” $\bm{x}_{m2}=$A {square or diamond with 4 edges, triangle with 3 edges}.”

### 27 Feb 2024 v2

@@ Abstract
Generative models can serve as surrogates for some real data sources by creating synthetic training datasets, but in doing so they may transfer biases to downstream tasks. We focus on protecting quality and diversity when generating synthetic training datasets. We propose quality-diversity generative sampling (QDGS), a framework for sampling data uniformly across a user-defined measure space, despite the data coming from a biased generator. QDGS is a model-agnostic framework that uses prompt guidance to optimize a quality objective across measures of diversity for synthetically generated data, without fine-tuning the generative model. Using balanced synthetic datasets generated by QDGS, we first debias classifiers trained on color-biased shape datasets as a proof-of-concept. By applying QDGS to facial data synthesis, we prompt for desired semantic concepts, such as skin tone and age, to create an intersectional dataset with a combined blend of visual features. Leveraging this balanced data for training classifiers improves fairness while maintaining accuracy on facial recognition benchmarks. Code available at: github.com/Cylumn/qd-generative-sampling.
@@ 1 Introduction
Balancing synthetic datasets is not a trivial task, since generative models harbor biases that relate to different semantic concepts. Some biases result from training deep generative models to replicate Internet data, where there are common issues of sampling selection bias. Internet data also contain social biases that result from disparities in language use among social groups, eliciting stereotypes and prejudices (Otterbacher 2018). Biases are intensified by selecting training approaches that amplify class and feature imbalances, such as some data augmentation and regularization methods (Balestriero, Bottou, and LeCun 2022). Thus, most existing generative models suffer from multiple biases that harm the creation of fair synthetic datasets. When creating these datasets for training downstream models, it is helpful to simultaneously consider multiple of these biases.
@@ 6 Results and Discussion
6.1 Spread of synthetic data distributions
@@
[…]
@@
Table 3: Ablation results: QDGS is helpful for either loss.
@@ B.2 Measure language prompts
  • $\bm{x}_{m_{1}}=$“{red, blue}.” $\bm{x}_{m_{2}}=$A {square, triangle}.”
  • $\bm{x}_{m_{1}}=$A {red, blue} shape.” $\bm{x}_{m_{2}}=$A {square or diamond with 4 edges, triangle with 3 edges}.”